1. a) A program has 200 instructions and runs in a processor in 5 ns. If each of the instructions take

10 clock cycles, what is the value of Clock Cycle Time?

b)A program has 125 instructions and total clock cycles consumed by the program is 31.25 ms.

What is the value of cock per instruction?


1
Expert's answer
2021-04-02T03:18:37-0400

Given,

Number of instruction = 200

Processor's running time = 5ns

Clock cycle = 10

Clock cycle time = 1/clock rate

"=\frac{5ns}{10}=0.5ns"

b)

CPU clock cycle = instruction count"\times" CPI

"\Rightarrow CPI=\frac{CPU \ clock \ cycle}{Instruction \ count}"


"\Rightarrow CPI = \frac{31.25ns}{125}ns"

"=0.25ns"